Fermi National Laboratory
 

Operation Summary for November 29, 2010 - December 1, 2010

 

Monday Nov. 29

The day shift began with Operations monitoring Tevatron (TeV) store 8315 (120.25E30) and the Recycler stash (175.2E10), with the Antiproton Source (Pbar) stacking (25.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Pbar personnel investigated issues with the Debuncher optical delay line. Mechanical personnel accessed MI-31 to remove equipment.

During the evening shift, Operations terminated store 8315 at 5 PM, put the TeV into HEP shot setup, and established store 8316 at 6:30 PM with an average initial luminosity (AIL) of 335E30. Operations reported that the TeV sector C4 vacuum controls VME communication failed.

Tuesday Nov. 30

The midnight shift began with Operations monitoring store 8316 (142.73E30) and the stash (133.3E10), with Pbar stacking (20.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported at 7:40 AM that store 8316 quenched at E2 due to instability in the cryo system transfer line. The Cryo Coordinator reported that the helium cold compressor B at CHL had failed midway through the shift.

During the day shift, TeV quench recovery continued. Operators prepared CDF, D-Zero, and the TeV A-E enclosures for access for routine maintenance and to repair the D11 power lead. Linac personnel switched from using the H- Source to the I- Source. Proton Source personnel spent the rest of the day tuning to restore normal intensity. All accesses had ended by 2 PM. Operations began the TeV turn-on and orbit checks. Pbar personnel began a 3.5-hour access into the rings to address Debuncher beam cooling issues and the 4-8 stochastic cooling; water technicians also found and repaired LCW leaks. Pbar resumed stacking at 1:45 PM.

During the evening shift, Operations put the TeV into shot setup at 6:20 PM and established store 8320 at 7:20 PM with an AIL of 334E30. The TeV antiproton longitudinal emittance increased, which caused a drop of 10E30 in luminosity shortly after the store began.

Wednesday Dec. 1

The midnight shift began with Operations monitoring store 8320 (152.64E30) and the stash (193.8E10), with Pbar stacking (16.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8320 at 7:30 AM and put the TeV into shot setup.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/01/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory